A bond with an annual coupon rate of 4.3% sells for $995. What is the bonds current yield? Round your answer to 2 decimal
Answer #1

Annual coupon=1000*4.3%=$43

Hence current yield=Annual coupon/Current value

=43/995

=4.32%(Approx).